# Inferring Ingrained Remote Information in AC Power Flows Using Neuromorphic Modality Regime

### v1

@@ III Performance Validation
The proposed methodology is validated through a co-simulation of Python and Matlab. The SNN model is built using Python and then integrated into the Simulink environment for simulations. The state of the inverter at Bus #6 in the IEEE 14-bus system in Fig. 3 is used for verification.
@@ III-A Coordination Control Results
In Fig. 4, we firstly illustrate the efficacy of reactive power sharing and average voltage regulation through two test cases: $\texttt{Case}_{1}$: a load at Bus #6 transitions from 0.45 pu to 0.9 pu; $\texttt{Case}_{2}$: a line outage occurs between Bus #9 and #14. In Fig. 4(a), the membrane potential $\bm{u}_{abc}$ are obtained from SNN, while the target values are represented as $\bm{m}_{abc}$. Before t = 0.5 s, the system maintains steady state, with $\bm{u}_{abc}$ remaining zero. However, with the onset of $\texttt{Case}_{1}$, $\bm{u}_{abc}$ are activated to match $\bm{m}_{abc}$, such that the spikes $\bm{spk}_{abc}$ are also generated. Since the membrane potential matches the target modulation signal in Fig. 4(a), the precision behind spike generation for the inverter at Bus #6 can be formally validated in Fig. 4(b). As evident from Fig. 4(c) and (d), the coordinated control outcomes for both $\texttt{Case}_{1}$ and $\texttt{Case}_{2}$ are satisfactorily achieved with the average voltage well regulated under several dynamic conditions.
@@ III-B Adaptability verification
To illustrate the adaptability of SNN to dynamic conditions, we employ the following cases: $\texttt{Case}_{3}$: utility grid voltage sag from 1 pu to 0.1 pu.; $\texttt{Case}_{4}$: line outage transpires between Bus #11 and #12. Conventionally, handling voltage sags necessitate a mode change between two control mechanisms [11]. However, in this work, this is carried out in seamless manner that adapts its operation only by sensing declining voltages locally. As depicted in Fig. 5(a) and (b), the system exhibits instability. However, upon integrating the FRT strategy, voltage and current are adaptively adjusted/safeguarded within the expected limits in response to the extent of the voltage sag, as illustrated in Fig. 5(c) and (d).
@@
Control parameters play a crucial role in system stability. Tuning a single set of parameters for a global set of scenarios is challenging. In system reconfiguration scenarios like a line outage in $\texttt{Case}_{4}$, the system can become unstable if the parameters are not adjusted properly, as illustrated in Fig. 6(a). Hence, we have integrated adaptive parameters into SNN, resulting in the improved stability of the system, as demonstrated in Fig. 6(b).
@@ III-C Energy Consumption Analysis
To assess the energy efficiency of SNN, we compare its average online operational energy consumption of 90 events with that of artificial neural networks (ANN) and binary-activated RNN, as presented in Table I. It is worth notifying that $P_{\mathrm{on}}$ and $P_{\mathrm{off}}$ denote the power consumption during event-on and event-off states, respectively. Similarly, $N_{\mathrm{on}}$ and $N_{\mathrm{off}}$ represent the count of active neurons in the event-on and event-off states, respectively with $E_{\mathrm{data}}$ denoting the energy consumption of each active neuron. Since binary RNN is also activated during spikes, $P_{\mathrm{off}}$ is 0 for SNN as compared to binary RNN that still remains active even in the absence of spikes. However under the presence of spikes, $P_{\mathrm{on}}$ of binary RNN and SNN are quite similar. In contrast, ANN continuously operates with real numbers, resulting in considerably higher energy consumption due to ADC stages for numeric values and the absence of synaptic plasticity property for memory leakage. Consequently, the energy efficiency advantage of SNN over ANN and binary RNN is quite evident.
@@
$P_{\mathrm{off}}$ [mW] $P_{\mathrm{on}}$ [mW] $N_{\mathrm{on}}$ $N_{\mathrm{off}}$ $E_{\mathrm{data}}$ [pJ]
@@
In Fig. 7(a), the accuracy of semantic sampling is demonstrated over the moving window based on the alignment of input and hidden spikes of SNN with respect to the sinusoidal signal $v_{a}[m]$. Since the inputs are sinusoidal signals, the internal spikes of SNN in the hidden layer are also in the same law aligned to the input signal as in Fig. 7(a). In Fig. 7(b), when the MSE in the moving window of $v_{d}$ from (1) is larger than threshold $\sigma_{th}^{V}$, the spikes are continuously generated as per the condition outlined in Algorithm I.
Figure 7: Spikes of hidden layer I in SNN deployed in inverter at Bus #6: (a) spikes generation alongside the input signal $v_{a}$, (b) spikes generation alongside the moving window.
@@ III-D Experimental Results
The performance of the proposed approach is validated experimentally in the setup shown in Fig. 8, which illustrates a two-bus islanded testbed. Two 7.5 kVA DC–AC converters (designated as Converters 1 and 2) operate with a bus voltage of 110 V are connected to each other supplying a resistive load via tie lines. The load resistance is 115 $\Omega$, with current sensing facilitating the objective of secondary controller. Since the spikes are generated directly in the controller and can not be directly sensed, its data is recorded by the Imperix Cockpit and re-plotted. The designed SNN is given by: $\bm{Layer}_{in,6\times 1}$, $\bm{Layer}_{hidden,256\times 2}$, $\bm{{Layer}}_{out,3\times 1}$}.
The results are shown in Fig. 9. In Fig. 9(a)-(d), we firstly present the results of load change. In Fig. 9(a), the load shifts from 115 $\Omega$ to 75 $\Omega$ at $t_{1}$. The output currents of converters $i_{oa1}$ and $i_{oa2}$ increase and are controlled to match. Due to the equivalent reactive load of the filter capacitor and line inductance, $v_{q1}$ is non-zero in Fig. 9(b). Considering the load in series with the line inductance and LC filter, resistive load changes cause both active and reactive power dynamics. Consequently, both $i_{d1}$ and $i_{q1}$ change, as illustrated in Fig. 9(c). The detection of $i_{d1}$, $i_{q1}$, $v_{d1}$, and $v_{q1}$ dynamics triggers events as per (4). Consequently, the spikes originating from the event-driven sampling are hereby generated. For a clear presentation of spikes, only the first 8 neurons are presented in the hidden layer. Finally, when the event is inactive, the SNN enters into an idle state.
In Fig. 9(e)-(h), we present the results of a voltage sag scenario. Converter 2’s output voltage is deliberately lowered to test the adaptive response of converter 1. In Fig. 9(e), converter 1’s output voltage $v_{a1}$ adeptly tracks $v_{a2}$ at $t_{2}$. This voltage matching results in controlled lower values for the two output currents, $i_{oa1}$ and $i_{oa2}$, without excessive overcurrent. The values of $v_{d1}$ and $v_{q1}$ in Fig. 9(f) shift closer to zero. Due to the voltage sag, both $i_{d1}$ and $i_{q1}$ decrease, as illustrated in Fig. 9(g). The detection of dynamics in $i_{d1}$, $i_{q1}$, $v_{d1}$, and $v_{q1}$ triggers the events, as shown in Fig. 9(h). The time scale of spikes aligns coherently with the detected event.
@@
Figure 9: Experimental results of load change and voltage sag using the proposed approach: (a) & (e) output voltages and currents of Converter 1 and 2 in phase A, (b) & (f) $v_{d}$ and $v_{q}$ of converter 1, (c) & (g) $i_{d}$ and $i_{q}$ of converter 1, (d) & (h) Event detection and spikes of the first 8 neurons in the hidden layer are presented.
@@ IV Conclusions and Future scope of work
This paper introduces a multi-modality regime using neuromorphic processing and semantics theory to represent power flow in a AC system having implicit information of remote DERs that can be typically used for cognition of power electronic dominated grids. This is made possible using SNNs deployed at each node not just as a means to communicate/coordinate between DERs, but also for enabling adaptive response during parameter mismatch and system reconfiguration issues. As opposed to the state-of-the-art grid-edge technologies, it offers the following advantages:
@@
    It dismisses the cyber layer, which prevents exogenous path arrivals for the attackers via the vulnerabilities.
@@
  4. 4.
    Its fast learning capabilities allow parameter adaptation for each DER based on the operational context (such as voltage sags and load changes) and can be effective in scenarios such as system expansion and reconfiguration.
As a future scope of work, we plan to explore the effectiveness of its powerful online learning capability against some of the unseen grid conditions and establish system resiliency without communications as an overarching layer. The complete scheme still needs to be further put into practice and validated with fully parallel memristor crossbar array circuits with improved qualitative realization of events.

### v2

@@ III Performance Validation
The proposed methodology is validated through a co-simulation of Python and Matlab. The SNN model is built using Python and then integrated into the Simulink environment for simulations. The state of the inverter at Bus #6 in the IEEE 14-bus system in Fig. 3 is used for verification. With reactive power sharing and average voltage regulation being the coordination objectives of the secondary controller for each inverter, fault ride-through (FRT) capability [16] is also seamlessly integrated into the SNN.
@@ III-A Coordination Control Results
In Fig. 4, we firstly illustrate the efficacy of reactive power sharing and average voltage regulation through two test cases: $\mathrm{\texttt{Case}_{1}}$: a load at Bus #6 transitions from 0.45 pu to 0.9 pu; $\mathrm{\texttt{Case}_{2}}$: a line outage occurs between Bus #9 and #14. In Fig. 4(a), the membrane potential $\mathrm{\mathbf{u}_{abc}}$ are obtained from SNN, while the target values are represented as $\mathrm{\mathbf{m}_{abc}}$. Before t = 0.5 s, the system maintains steady state, with $\mathrm{\mathbf{u}_{abc}}$ remaining zero. However, with the onset of $\mathrm{\texttt{Case}_{1}}$, $\mathrm{\mathbf{u}_{abc}}$ are activated to match $\mathrm{\mathbf{m}_{abc}}$, consequently leading to generation of spikes $\mathrm{\mathbf{spk}_{abc}}$. Since the membrane potential matches the target modulation signal in Fig. 4(a), the precision behind spike generation for the inverter at Bus #6 can be formally validated in Fig. 4(b). As evident from Fig. 4(c) and (d), the coordinated control outcomes for both $\mathrm{\texttt{Case}_{1}}$ and $\mathrm{\texttt{Case}_{2}}$ are satisfactorily achieved with the average voltage well regulated under several dynamic conditions.
@@ III-B Adaptability verification
To illustrate the adaptability of SNN to dynamic conditions, we employ the following cases: $\mathrm{\texttt{Case}_{3}}$: utility grid voltage sag from 1 pu to 0.1 pu.; $\mathrm{\texttt{Case}_{4}}$: line outage transpires between Bus #11 and #12. Conventionally, handling voltage sags necessitate a mode change between two control mechanisms [16]. However, in this work, this is carried out in seamless manner that adapts its operation only by sensing declining voltages locally. As depicted in Fig. 5(a) and (b), the system exhibits instability. However, upon integrating the FRT strategy, voltage and current are adaptively adjusted/safeguarded within the expected limits in response to the extent of the voltage sag, as illustrated in Fig. 5(c) and (d).
@@
Control parameters play a crucial role in system stability. Tuning a single set of parameters for a global set of scenarios is challenging. In system reconfiguration scenarios such as the line outage in $\mathrm{\texttt{Case}_{4}}$, the system can become unstable if the parameters are not adjusted properly, as illustrated in Fig. 6(a). Hence, adaptation of SNN weights using power flow dynamics modifies SNN parameters, resulting in the improved stability of the system, as demonstrated in Fig. 6(b).
@@ III-C Energy Consumption Analysis
To assess the energy efficiency of SNN, we compare its average online operational energy consumption of 90 events with that of artificial neural networks (ANN) and binary-activated recurrent neural networks (RNN), as presented in Table I. It is worth notifying that $\mathrm{P_{\mathrm{on}}}$ and $\mathrm{P_{\mathrm{off}}}$ denote the power consumption during event-on and event-off states, respectively. Similarly, $\mathrm{N_{\mathrm{on}}}$ and $\mathrm{N_{\mathrm{off}}}$ represent the count of active neurons in the event-on and event-off states, respectively with $\mathrm{E_{\mathrm{data}}}$ denoting the energy consumption of each active neuron. Even though the event-driven operation of binary RNN and SNN are quite similar, they basically differ in their off-time power consumption (in the absence of spikes), where $\mathrm{P_{\mathrm{off}}}$ is 0 for SNN as compared to 6.06 mW for binary RNN. However under the presence of spikes, $\mathrm{P_{\mathrm{on}}}$ of binary RNN and SNN are quite similar. As opposed to binary spikes based data processed asynchronously in SNN and binary RNN, ANN operates synchronously to process floating point numbers based data, resulting in considerably higher energy consumption due to ADC stages for numeric values. Consequently, the energy efficiency advantage of SNN over ANN and binary RNN is quite evident.
@@
$\mathrm{P_{\mathrm{off}}}$ [mW] $\mathrm{P_{\mathrm{on}}}$ [mW] $\mathrm{N_{\mathrm{on}}}$ $\mathrm{N_{\mathrm{off}}}$ $\mathrm{E_{\mathrm{data}}}$ [pJ]
@@
In Fig. 7(a), the accuracy of asynchronous event-driven sampling is demonstrated over the moving window based on the alignment of input and hidden spikes of SNN with respect to the sinusoidal signal $\mathrm{v_{a}[m]}$. Since the inputs are sinusoidal signals, the spikes collected in the hidden layer of the SNN are aligned to the input signal as in Fig. 7(a). In Fig. 7(b), when the MSE in the moving window of $\mathrm{v_{d}}$ from (1) is larger than threshold $\mathrm{\sigma_{th}^{V}}$, spikes are continuously generated as per the condition outlined in Algorithm I.
Figure 7: Spikes of hidden layer I in SNN deployed in inverter at Bus #6: (a) spikes generation alongside the input signal $\mathrm{v_{a}}$, (b) spikes generation alongside the moving window.
@@ III-D Experimental Results
The performance of the proposed approach is validated experimentally in the setup shown in Fig. 8, which illustrates a two-bus islanded testbed. Two 7.5 kVA DC–AC converters (designated as Converters 1 and 2) operating with a reference bus voltage of 110 V are connected to each other via tie-lines to supply a resistive load. The load resistance is 115 $\mathrm{\Omega}$. Since the spikes are generated inside the controller and can not be directly sensed, we recorded the spikes-based data using the Imperix Cockpit and re-plotted it in Fig. 9(d) and 9(h). The designed SNN is given by: $\mathrm{\mathbf{Layer}_{in,6\times 1}}$, $\mathrm{\mathbf{Layer}_{hidden,256\times 2}}$, $\mathrm{\mathbf{{Layer}}_{out,3\times 1}}$}.
The results are shown in Fig. 9. In Fig. 9(a)-(d), we firstly present the results of load change. In Fig. 9(a), the load shifts from 115 $\mathrm{\Omega}$ to 75 $\mathrm{\Omega}$ at $\mathrm{t_{1}}$. The output currents of converters $\mathrm{i_{oa1}}$ and $\mathrm{i_{oa2}}$ increase and are controlled to match. Due to the equivalent reactive load of the filter capacitor and line inductance, $\mathrm{v_{q1}}$ is non-zero in Fig. 9(b). Considering the load in series with the line inductance and LC filter, resistive load changes cause both active and reactive power dynamics. Consequently, both $\mathrm{i_{d1}}$ and $\mathrm{i_{q1}}$ change, as illustrated in Fig. 9(c). The detection of $\mathrm{i_{d1}}$, $\mathrm{i_{q1}}$, $\mathrm{v_{d1}}$, and $\mathrm{v_{q1}}$ dynamics triggers events as per (4). Consequently, the spikes originating from the event-driven sampling are hereby generated. For a clear presentation of spikes, only the first 8 neurons are presented in the hidden layer. Finally, when the event is inactive, the SNN enters into an idle state.
In Fig. 9(e)-(h), we present the results of a voltage sag scenario. Converter 2’s output voltage is deliberately lowered to test the adaptive response of converter 1. In Fig. 9(e), converter 1’s output voltage $\mathrm{v_{a1}}$ adeptly tracks $\mathrm{v_{a2}}$ at $\mathrm{t_{2}}$. This voltage matching results in controlled lower values for the two output currents, $\mathrm{i_{oa1}}$ and $\mathrm{i_{oa2}}$, without excessive overcurrent. The values of $\mathrm{v_{d1}}$ and $\mathrm{v_{q1}}$ in Fig. 9(f) shift closer to zero. Due to the voltage sag, both $\mathrm{i_{d1}}$ and $\mathrm{i_{q1}}$ decrease, as illustrated in Fig. 9(g). The detection of dynamics in $\mathrm{i_{d1}}$, $\mathrm{i_{q1}}$, $\mathrm{v_{d1}}$, and $\mathrm{v_{q1}}$ triggers the events, as shown in Fig. 9(h). The time scale of spikes aligns coherently with the detected event.
@@
Figure 9: Experimental results of load change and voltage sag using the proposed approach: (a) & (e) output voltages and currents of Converter 1 and 2 in phase A, (b) & (f) $\mathrm{v_{d}}$ and $\mathrm{v_{q}}$ of converter 1, (c) & (g) $\mathrm{i_{d}}$ and $\mathrm{i_{q}}$ of converter 1, (d) & (h) Event detection and spikes of the first 8 neurons in the hidden layer are presented.
@@ IV Conclusions and Future scope of work
This paper leverages neuromorphic computing principles using a SNN at each grid-edge power electronic converter that enables them to comprehend grid disturbances by observing local power dynamics across the tie-lines in an asynchronous, event-driven manner and provide online adaptive response to it accordingly. As opposed to the state-of-the-art grid-edge technologies, it offers the following advantages:
@@
    It dismisses coordination via cyber layer, thereby preventing exogenous path arrivals for cyber attackers and further security concerns.
@@
As a future scope of work, we plan to explore the effectiveness of its powerful online learning capability against some of the unseen grid conditions and establish system resiliency without communications as an overarching layer.